Determine whether the graphs of the equations are perpendicular.
7x-6y=5
7y-6x=6
Are the graphs of the given equations perpendicular
yes or no. Show your work.

Answer by stanbon(75887)

Determine whether the graphs of the equations are perpendicular.
7x-6y=5
7y-6x=6
-------
Solve both for "y":
y = (7/6)x - (5/6)
y = (7/6)x - 1
----
Slopes are the same and intercepts different so the
line are not identical but they are parallel.
Cheers,
Stan H.
-----------
Are the graphs of the given equations perpendicular
yes or no. Show your work.